OpenStudy (anonymous):

Gunther drew a circle. The radius of the circle is 20 inches. He uses the formula below to determine the area of the circle. A="pi"radius to the power of two What is the area in square inches of Gunther's circle? Leave your answers in terms of "pi""

OpenStudy (anonymous):

A=pi*r^2= pi*20^2=40pi square inches

OpenStudy (anonymous):

thanks